Shift report 20030225 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 25 2003 - 16:21:00 EST

  • Next message: Flemming Videbaek: "Wednesday access"
    ------------------------------------------------------------
    Date: Tue Feb 25 16:20:59 2003, Shift: 08:00-16:00
    Supervisor: BHS, with: 
    Log book#14 pages 36 - 36
    
    Shift Summary:
    ------------------------------------------------------------
    1: Machine has had a beautiful store all shift, but because of HV problem we have not taken any physics data. 
    
    2: FV has player a lot with the triggers. The full FS trigger (FFS+H2) is now cabled in, plus other transparent changes - see elog for details. Fv also fixed a problem with an H2 channel.
    
    3: D5 tripped once, but came back on nicely. We noted that T5 draws a very high current on cathode 2 - see elog. Expert?
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 25 2003 - 16:22:19 EST